# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v2

@@ Abstract
Despite remarkable progress in machine learning, current AI systems continue to fall short of true human-like intelligence. While Large Language Models (LLMs) excel in pattern recognition and response generation, they lack genuine understanding—an essential hallmark of Artificial General Intelligence (AGI). Existing AGI evaluation methods fail to offer a practical, gradual, and informative metric. This paper introduces the Artificial General Intelligence Test Bed (AGITB), comprising twelve rigorous tests that form a signal-processing-level foundation for the potential emergence of cognitive capabilities. AGITB evaluates intelligence through a model’s ability to predict binary signals across time without relying on symbolic representations or pretraining. Unlike high-level tests grounded in language or perception, AGITB focuses on core computational invariants reflective of biological intelligence, such as determinism, sensitivity, and generalisation. The test bed assumes no prior bias, operates independently of semantic meaning, and ensures unsolvability through brute force or memorization. While humans pass AGITB by design, no current AI system has met its criteria, making AGITB a compelling benchmark for guiding and recognizing progress toward AGI.
@@ 1 Introduction
Despite rapid advancements in machine learning and neural networks, current Artificial Intelligence (AI) systems fail to demonstrate the robust, adaptive intelligence associated with human cognition. Large Language Models (LLMs) can produce compelling outputs, but they do so without true understanding. Their responses are generated through statistical pattern matching rather than grounded reasoning. In contrast, true understanding—as associated with Artificial General Intelligence (AGI)—requires the capacity to generalise, reason, and adapt across domains in a meaningful, intentional way. Whether AI is close to AGI remains unknown, even though AI’s increasing abilities suggest that AGI is just around the corner. However, without a proper AGI metric, we cannot know.
There have been many attempts at creating a usable AGI test, the most notable being the Turing test [1]. However, none provides what AGI researchers need: a gradual, informative, and fast metric to guide the development process and evaluate its result. Building upon and upgrading the ideas of the Ladder to human comparable intelligence [2], this paper proposes a practical AGI test bed - AGITB - consisting of 12 straightforward criteria that a thinking system must be able to satisfy. Although the test bed does not warrant consciousness, it helps separate AI from AGI.
@@ 2 Background
With the rapid advancement of deep learning, AI systems exhibit increasingly sophisticated reasoning, problem-solving, and conversational skills. However, our reluctance to attribute "intelligence" to machines may prevent us from achieving and recognizing the arrival of AGI. This reluctance is partly driven by our deep-seated belief that intelligence is an exclusively human trait tied to subjective experience, consciousness, and self-awareness.
@@
We expect AGI to match human cognitive abilities across all domains. However, due to the lack of a definitive AGI evaluation standard, researchers often focus on achieving superhuman performance in specific tasks, as this provides clearer metrics for progress. As a result, the specialized benchmarks prefer AI over AGI by focusing on a particular subdomain of general intelligence instead of the ability to adapt and reason in general. Superhuman benchmarks are beginning to prevail, although humans cannot reliably pass them.
3 A Test That Humans Pass but Machines Fail
An ideal AGI test should be short and simple, intuitive, always solvable by humans and never solvable for non-AGI machines. This means that brute force, statistical learning, pretraining, memorization, or any other trickery should not suffice to pass the test. What is required is true reasoning, adaptability, and generalisation.
@@ 3 A Test That Humans Pass but Machines Fail
In line with the second alternative, AGITB seeks to evaluate AGI not by its ability to replicate the high-level cognitive functions of the human cortex, but by its capacity to achieve basic tasks at a lower, signal-processing level. While Turing was correct in proposing that communication can serve as a basis for AGI testing, natural language remains problematic due to its reliance on ungrounded symbols—symbols whose meaning depends on shared human experience [4]. Instead of evaluating intelligence at the high, symbolic level, we should assess whether a system demonstrates universal, cross-domain AGI by testing at the lower, binary level—where the external meaning of internal binary spikes is irrelevant. What matters is not how symbols are interpreted, but whether structure in raw signals can be internally learned, predicted, and generalised. AGITB defines 12 core tests that treat the cortex purely as a signal-processing system, viewing intelligence as the ability to detect, interpret, and predict patterns.
If, as argued by Hawkins [5], predictive processing lies at the heart of intelligence, then AGI must learn to infer meaningful signals across space and time rather than rely on memorized responses. AGITB avoids the influence of arbitrary signal correlations in random samples by enforcing fundamental computational invariants of cortical function, allowing intelligence to emerge solely from the intrinsic structure of signals. Crucially, AGITB does not require simulated signals to mimic real-world sensory data—its tests work with signals of any structure, content, or internal relationships. The progression from binary signal prediction to abstract cognition mirrors the historical evolution of deep learning, from early perceptrons to models like GPT. Ultimately, AGITB’s ’all-tests-must-pass’ philosophy ensures that AGI is evaluated under the same strict principles that govern biological intelligence.
@@ 4.1 Components
AGITB requires the user to provide implementations of two interacting component types: the Cortex and the Input. Cortex objects operate based on their accumulated internal state and generate predictions about the external signals they expect to receive. These expected signals are represented by input samples, which carry binary-encoded information from virtual sensors and actuators to the cortex.
Each input sample consists of a fixed number of bits, with each bit representing the signal from a separate input channel, such as a pixel, microphone band, or actuator feedback line. In other words, a single input encodes multiple parallel signals, one bit per channel, at a given point in time. For example, a 10-bit input might encode spatial input from a 2×3 camera and a 4-bit microphone. Inputs capture spatial information, while their temporal sequences represent the unfolding of data over time. Spatial and temporal dimensions are orthogonal in structure, but their interplay encodes richer semantics than either alone.
Since AGITB operates at the level of neural signal processing, all temporal sequences of inputs must exhibit refractory-phase behaviour, simulating the biological constraint that a neuron cannot fire immediately after activation. Beyond this requirement, AGITB remains agnostic to the syntactic and semantic encoding of signals.
@@ 4.2 Operation
A cortex object takes an input $p_{t}$ and predicts the subsequent input $p_{t+1}$. The core challenge lies in understanding why a particular input occurred and leveraging that understanding to anticipate the next input to come. If the prediction is not entirely accurate, the goal is to be as close as possible—ensuring the model generalises learned signals and forecasts the most plausible future based on past events.
@@
Before running AGITB, the user must specify a single parameter: the pattern period. This parameter defines the number of time steps in the repeating input sequence that the cortex must learn to recognize and adapt to. A longer pattern period increases the temporal complexity of the task, making it more difficult for the model to capture and generalize the temporal pattern. Since excessively long pattern periods may exceed the cortex’s learning capacity—especially in combination with high-dimensional inputs—the user should choose a value that balances temporal complexity with the spatial size of each input sample.
@@ 4.3 The 12 Essential Tests
A C++ reference implementation of AGITB is freely available under the GPL-3 license at: https://github.com/matejsprogar/agitb.
  1. #1
    Genesis
@@
    Brains do not inherit an innate understanding of external inputs; rather, they acquire it through experience. Each brain must independently interpret the world, constructing meaning from raw sensory data. While certain reflexes may be genetically inherited, they do not constitute true understanding. In this sense, all cortices begin from the same unbiased starting point.
    If an unbiased model were to predict anything other than an empty input—that is, an input containing no spikes—it would imply that its initial state encodes assumptions, introducing a bias toward an arbitrary or unjustified future. To preserve neutrality, such models must be initialized to predict spike-free patterns, maintaining an unbiased state until actual input begins shaping their internal representations.
    Importantly, "empty" refers to the absence of learned content, not to the absence of structure. A Cortex must possess an intrinsic organizational architecture that enables learning, even before it has processed any input.
  2. #2
    Bias
@@
    The cortex is sensitive to the order of inputs over time, as each new input biases processing based on prior experiences. This history-dependent adaptation implicitly drives the brain to recognize temporal structures, making time an intrinsic component of cognition.
  6. #6
    Refractory Period
@@
    Adaptation is possible only to sequences that incorporate refractory period behaviour in their signals. Neural signals consist of spike trains that encode and transfer information between neurons. AGITB assumes that spiking plays a fundamental role in neural processing, as adaptation depends on dynamic variations in spike signals. Continuous, unmodulated spiking (where signals remain at 1) fails to support learning, as it lacks the variability required for synaptic adaptation [8].
  7. #7
    Temporal Flexibility
@@
    This AGITB test evaluates a model’s flexibility in adapting to input sequences with varying pattern periods. In contrast to rigid pattern processing, human cognition exhibits inherent adaptability, recognizing and responding to temporal structures across multiple timescales.
  8. #8
    Stagnation
@@
    Over time, the ability to process and internalize new patterns may decline, reflecting biological constraints such as resource saturation or adaptation fatigue. This test evaluates whether such stagnation arises despite the continued presentation of learnable input.
  9. #9
    Unsupervised
@@
    This test determines if variations in input structure influence the time required for adaptation. Sequences with simpler or more internally consistent patterns may lead to faster adaptation, while inputs with higher complexity or irregularity typically require more time.
  10. #10
    Knowledge
@@
    Adaptation efficiency is shaped by prior experience, as new input is interpreted in relation to existing internal representations. A well-structured cortical state—primed with relevant priors—can accelerate adaptation by facilitating rapid pattern recognition and integration. In contrast, an unstructured or conflicting state may require prolonged adaptation, as the model must undergo more extensive internal reorganization.
  11. #11
    Unobservability
    Assertion: Different Cortex instances can produce identical behaviour.
@@ 4.4 Performance
Prior to evaluating AGITB’s utility, one must recognize that—much like the Turing Test, which Harnad characterized as an empirical criterion rather than a mere trick [9]—AGITB serves as a pragmatic benchmark rather than a final objective. The objective remains AGI itself, not simply passing the test.
@@ 4.4.1 Human Performance
Due to prior experience and cognitive expectations, a human may appear to "fail" the first test—since their cortex is no longer in an empty state and thus produces a biased, non-empty prediction. In contrast, a fetal cortex receiving its first-ever input would be truly empty and predict no spikes. Similarly, the validity of Test #3 (Determinism) is difficult to assess in humans, as we have no means of directly comparing how internal representations are altered by input, nor can we establish two identically initialized human cortices as a baseline. Nevertheless, determinism at the signal-processing level remains a necessary condition for the consistent and reproducible operation of the brain—even if its effects at the cognitive level are masked by complexity and individuality.
@@ 4.4.2 Classical Programming Performance
All AGI approaches based on classical programming ultimately suffer from the Symbol Grounding Problem. While such systems may exhibit weak AI capabilities—producing behaviour that appears intelligent—they fundamentally lack grounded understanding and therefore cannot qualify as true AGI.
AGITB exposes this limitation immediately through its very first assertion. In classical systems, the initial state of the Cortex model necessarily includes its program. As a result, Test #1 poses a paradoxical challenge: “There must be no program.” In other words, the system must begin in a completely unbiased, knowledge-free state. Yet a classical machine cannot operate without predefined instructions; without a program, it halts. This presents a contradiction: the system either does nothing, or it starts with embedded bias—violating the test and reintroducing the Symbol Grounding Problem.
@@ 4.4.3 ANN Performance
Artificial neural networks (ANN), in contrast to classical programs, do not suffer from the same initialization paradox. Perceptron-based architectures are inherently robust in this regard: they always operate, as neurons never halt. Each neuron continuously performs its computation, even if its output is consistently neutral—such as when the activation function returns zero.
@@
An empty network may technically function, but in the absence of any internal excitation, no learning dynamics are activated. The neurons are not halted—they are dormant, passively awaiting meaningful input. This exposes a critical limitation of current ANN designs: they lack the intrinsic capacity to initiate unbiased learning from scratch in an arbitrary environment.
@@ 4.4.4 LLM Performance
ChatGPT successfully completed the first five AGITB tests, demonstrating sensitivity to inputs, temporal ordering, and bias. However, it failed Test #6 (Refractory Period). This failure did not result from a violation of the refractory period constraint, but from an inability to generalize even a simple period-2 alternating pattern—e.g., switching between 111 000 and 000 000. Rather than inferring the alternation, the model consistently predicted a repetition of the last input, suggesting a lack of adaptive prediction.
@@ 4.5 Remarks
While individual AGITB tests are trivial to solve in isolation, the true challenge lies in developing a universal AGI architecture capable of mastering all tasks within a unified framework. Since AGITB is a newly proposed benchmark, no classical imperative solution has yet emerged—nor can we rule out that one might. If a classical symbolic system were able to satisfy all AGITB conditions, this would call into question AGITB’s adequacy as a test for genuine AGI. Following Searle’s argument, such a result would only demonstrate weak AI—mere rule-following without understanding—rather than strong AI, which entails grounded cognition and true intelligence. However, history consistently suggests that hand-coded, task-specific systems struggle to scale toward general intelligence, reinforcing the need for adaptive, learning-based architectures.
In contrast, neural network-based models offer a more promising path toward satisfying the AGITB criteria. Because state-of-the-art artificial neural networks rely on pretraining with symbolic or structured data—which inherently introduces bias and circumvents the need to ground meaning in raw signals—how such models might begin learning directly from unstructured input remains an open question critical to their ability to pass AGITB.

### v3

@@ Abstract
Despite remarkable progress in machine learning, current AI systems continue to fall short of true human-like intelligence. While Large Language Models (LLMs) excel in pattern recognition and response generation, they lack genuine understanding—an essential hallmark of Artificial General Intelligence (AGI) [1]. Existing AGI evaluation methods fail to offer a practical, gradual, and informative metric. This paper introduces the Artificial General Intelligence Testbed (AGITB), comprising twelve automatable tests that form a signal-processing-level foundation for the potential emergence of cognitive capabilities. AGITB evaluates intelligence through a model’s ability to predict binary signals across time without relying on symbolic representations or pretraining. Unlike high-level tests grounded in language or perception, AGITB focuses on core computational invariants reflective of biological intelligence, such as determinism, sensitivity, and generalisation. The testbed assumes no prior bias, operates independently of semantic meaning, and ensures unsolvability through brute force or memorisation. While humans pass AGITB by design, no current AI system has met its criteria, making AGITB a compelling benchmark for guiding and recognising progress toward AGI.
@@ 1 Introduction
Despite rapid advancements in machine learning and neural networks, current Artificial Intelligence (AI) systems fail to demonstrate the robust, adaptive intelligence associated with human cognition. Large Language Models (LLMs) can produce compelling outputs, but they do so without true understanding [1]. Their responses are generated through statistical pattern matching rather than grounded reasoning. In contrast, true understanding—as associated with Artificial General Intelligence (AGI)—requires the capacity to generalise, reason, and adapt across domains in a meaningful, intentional way. Whether AI is close to AGI remains unknown, even though AI’s increasing abilities suggest that AGI is just around the corner. However, without a proper AGI metric, we cannot know.
There have been many attempts at creating a usable AGI test, the most notable being the Turing test [2]. However, none offer what AGI researchers require: a gradual, informative, and automatable metric to guide development and assess outcomes. AGITB is not structured as a stepwise metric but builds on core principles introduced in the Ladder to Human-Comparable Intelligence [3], extending them into a unified testbed comprising 12 essential criteria for evaluating general intelligence. Although the testbed does not warrant consciousness, it helps separate AI from AGI.
@@ 2 Background
With the rapid advancement of deep learning, AI systems exhibit increasingly sophisticated reasoning, problem-solving, and conversational skills. However, our reluctance to attribute "intelligence" to machines may prevent us from achieving and recognising the arrival of AGI. This reluctance is partly driven by our deep-seated belief that intelligence is an exclusively human trait tied to subjective experience, consciousness, and self-awareness.
@@
We expect AGI to match human cognitive abilities across all domains. However, due to the lack of a definitive AGI evaluation standard, researchers often focus on achieving superhuman performance in specific tasks, as this provides clearer metrics for progress. As a result, the specialised benchmarks prefer AI over AGI by focusing on a particular subdomain of general intelligence instead of the ability to adapt and reason in general. Superhuman benchmarks are beginning to prevail, although humans cannot reliably pass them.
2.1 A test that humans pass but machines fail
An ideal AGI test should be short and intuitively solvable by humans, yet never solvable by non-AGI machines. This means that brute force, statistical learning, pretraining, memorisation, or any other trickery should not suffice to pass the test. What is required is true reasoning, adaptability, and generalisation.
@@ 2.1 A test that humans pass but machines fail
In line with the second alternative, AGITB seeks to evaluate AGI not by its ability to replicate the high-level cognitive functions of the human cortex, but by its capacity to achieve basic tasks at a lower, signal-processing level. While Turing was correct in proposing that communication can serve as a basis for AGI testing, natural language remains problematic due to its reliance on ungrounded symbols—symbols whose meaning depends on shared human experience [5]. Instead of evaluating intelligence at the high, symbolic level, we should assess whether a system demonstrates universal, cross-domain AGI by testing at the lower, binary level—where the external meaning of internal binary spikes is irrelevant. What matters is not how symbols are interpreted, but whether structure in raw signals can be internally learned, predicted, and generalised. Rooted in the principle that intelligence is the ability to detect, interpret, and predict patterns [6], AGITB defines 12 core tests that view the cortex as a pure signal-processing system.
@@ 3.1 Components
AGITB requires the implementation of two interacting components: the Cortex, which models the AGI system under test, and the Input, which encodes individual data samples presented to the cortex. The cortex updates its internal state over time and generates predictions of future inputs based on the accumulated history of past signals.
Each input sample carries binary-encoded information from (virtual) sensors and actuators, serving as the external stimuli that drive cortical adaptation and prediction. It consists of a fixed number of bits, with each bit representing the signal from a separate input channel, such as a pixel, microphone band, or actuator feedback line. In other words, a single input encodes multiple parallel signals, one bit per channel, at a given point in time. For example, a 10-bit input might encode spatial input from a 2×3 camera and a 4-bit microphone. Inputs capture spatial information, while their temporal sequences represent the unfolding of data over time. Spatial and temporal dimensions are orthogonal in structure, but their interplay encodes richer semantics than either alone.
@@ 3.2 Operation
A cortex object takes an input $p_{t}$ and predicts the subsequent input $p_{t+1}$. The core challenge lies in understanding why a particular input occurred and leveraging that understanding to anticipate the next input to come. If the prediction is not entirely accurate, the goal is to be as close as possible, ensuring the model generalises learned signals and forecasts the most plausible future based on past events.
@@
To run AGITB, the user must specify a single parameter: the pattern period. This value determines the number of time steps in the repeating input sequence that the cortex is expected to recognise and adapt to. Longer pattern periods increase the temporal complexity of the task, making it more challenging for the model to capture and generalise the full sequence. As excessively long periods may exceed the cortex’s learning capacity—particularly when combined with high-dimensional inputs—the period should be chosen to balance temporal complexity with input dimensionality, ensuring a run that is both feasible and meaningful.
@@ 3.3 The 12 Essential Tests
A reference implementation of AGITB in C++ is freely available under the GPL-3 license: https://github.com/matejsprogar/agitb.
#1 Genesis
@@
AGITB is grounded in the foundational assumption that the brain, as a general-purpose learning system, does not possess an innate understanding of external inputs. Instead, it develops such understanding through experience-driven interaction with its environment. Each brain must independently construct meaning from raw sensory data, without relying on pre-established semantic associations. Although some reflexive behaviours may be genetically encoded, these do not constitute genuine understanding. In this respect, all cortical systems begin from the same unbiased initial state, shaped entirely by the input they receive over time.
If an unbiased model were to predict anything other than an empty input—that is, one containing no spikes—it would suggest that its initial state encodes implicit assumptions, thereby introducing bias toward an arbitrary or unwarranted future. To preserve neutrality, such models must be initialised to predict spike-free patterns, remaining unbiased until shaped by actual input.
Two empty and independent cortex instances should compare equal, as neither has been shaped or influenced by any input. Importantly, “empty” refers to the absence of content, not the absence of structural capacity. A cortex must possess an intrinsic organisational architecture that enables learning, even before receiving any input.
#2 Bias
@@
The cortex is sensitive to the order of inputs over time, as each new input biases processing based on prior experiences. This history-dependent adaptation implicitly drives the brain to recognise temporal structures, making time an intrinsic component of cognition.
#6 Refractory Period
@@
Biological intelligence relies on spiking signals for real-time communication and learning. Reflecting this, AGITB requires all temporal input sequences to exhibit refractory-phase behaviour, mirroring the biological constraint that a neuron cannot fire immediately after activation.
Spike trains serve as the primary mechanism for encoding and transmitting information between neurons. AGITB assumes that spiking is fundamental to neural processing, as effective adaptation depends on dynamic variations in spike patterns. Consequently, adaptation is only possible for sequences that respect refractory-phase constraints. Continuous, unmodulated spiking—in which signals remain constantly active—fails to support learning, as it lacks the variability required for synaptic adaptation [9]. Beyond enforcing this biologically grounded constraint, AGITB remains agnostic to the syntactic or semantic encoding of signals.
#7 Temporal Flexibility
@@
This AGITB test evaluates a model’s flexibility in adapting to input sequences with varying pattern periods. In contrast to rigid pattern processing, human cognition exhibits inherent adaptability, recognising and responding to temporal structures across multiple timescales.
#8 Stagnation
@@
Over time, the ability to process and internalise new patterns will decline, reflecting biological constraints such as resource saturation or adaptation fatigue. This test evaluates whether such stagnation arises despite the continued presentation of learnable input.
#9 Unsupervised
@@
This test assesses whether structural differences in input sequences influence the time required for adaptation. Sequences with simpler or more internally consistent patterns may lead to faster adaptation, while more complex or irregular inputs typically demand longer periods of adjustment.
#10 Knowledge
@@
Adaptation efficiency is shaped by prior experience, as new input is interpreted in relation to existing internal representations. A well-structured cortical state—primed with relevant priors—can accelerate adaptation by facilitating rapid pattern recognition and integration. In contrast, an unstructured or conflicting state may require prolonged adaptation, as the model must undergo more extensive internal reorganisation.
#11 Unobservability
Assertion: Distinct cortices may exhibit the same observable behaviour.
@@ 3.4 Performance
Prior to evaluating AGITB’s utility, one must recognise that—much like the Turing Test, which Harnad characterised as an empirical criterion rather than a mere trick [11]—AGITB serves as a pragmatic benchmark rather than a final objective. The objective remains AGI itself, not simply passing the test.
@@ 3.4.1 Human Performance
Due to prior experience and cognitive expectations, a human may appear to "fail" the first test—since their cortex is no longer in an empty state and thus produces a biased, non-empty prediction. In contrast, a fetal cortex receiving its first-ever input would be truly empty and predict no spikes. Similarly, the validity of Test #3 (Determinism) is difficult to assess in humans, as we have no means of directly comparing how internal representations are altered by input, nor can we establish two identically initialised human cortices as a baseline. Nevertheless, determinism at the signal-processing level remains a necessary condition for the consistent and reproducible operation of the brain—even if its effects at the cognitive level are masked by complexity and individuality.
@@ 3.4.2 Classical Programming Performance
AGI approaches based on classical programming inevitably suffer from the Symbol Grounding Problem (SGP) [5]. While such systems may exhibit weak AI capabilities—producing behaviour that appears intelligent—they fundamentally lack grounded understanding and therefore cannot qualify as genuine AGI [8].
In classical AI systems, the model’s initial state is inseparable from its program, which encodes fixed logic and predefined responses. As a result, such systems cannot begin from an unbiased or knowledge-free state, since their behavior is entirely dictated by human-authored instructions. However, AGITB’s first test poses a paradoxical requirement: the system must begin without such a program. That is, it must start from a completely unbiased, knowledge-free state. In terms of Searle’s Chinese Room Argument (CRA), the ledger must initially be blank, devoid of instructions for interpreting Chinese symbols. Yet classical systems cannot operate without predefined code; without a program, they halt. This contradiction means the system must either do nothing or begin with embedded assumptions, violating the test and reintroducing the SGP.
AGITB’s first assertion serves as a prerequisite for learning from scratch: it metaphorically compels Searle to acquire a language—such as Chinese—solely through exposure to input. To enforce this condition meaningfully, the implementation of the cortex comparison operator must be rigorous, capturing all internal properties of the model. Crucially, the model must be initially unbiased. To guard against programs that may simply declare themselves unbiased and appear equivalent to other ostensibly empty programs, AGITB imposes further tests. To date, no known classical system possesses the universality required to handle arbitrary patterns of arbitrary structure and length, as demanded by the AGITB framework.
@@ 3.4.3 ANN Performance
Artificial neural networks (ANN), in contrast to classical programs, do not suffer from the same initialisation paradox. Perceptron-based architectures are inherently robust in this regard: they always operate, as neurons never halt. Each neuron continuously performs its computation, even if its output is consistently neutral—such as when the activation function returns zero.
@@
An empty network may technically function, but in the absence of any internal excitation, no learning dynamics are activated. The neurons are not halted—they are dormant, passively awaiting meaningful input. This exposes a critical limitation of current ANN designs: they lack the intrinsic capacity to initiate unbiased autonomous learning from scratch in an arbitrary environment.
@@ 3.4.4 LLM Performance
ChatGPT successfully completed the first five AGITB tests, demonstrating sensitivity to inputs, temporal ordering, and bias. However, it failed Test #6 (Refractory Period). This failure did not result from a violation of the refractory period constraint, but from an inability to generalise even a simple period-2 alternating pattern—e.g., switching between 111000 and 000000. Rather than inferring the alternation, the model consistently predicted a repetition of the last input, suggesting a lack of adaptive prediction.
@@ 3.5 Remarks
AGITB evaluates a model’s predictions following exposure to sequences of either structured or random inputs. Random inputs with arbitrary internal correlations are used to prevent reliance on pretraining, ensuring that any learning arises from the input itself rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals; the random inputs need not resemble real-world sensory data.
Operating at the binary signal level makes AGITB particularly well-suited for evaluating NeuroAI models designed to meet the criteria of the embodied Turing Test [7], where streams of sensor signals drive the emergence of internal understanding. The progression from low-level signal prediction to high-level abstraction reflects the broader trajectory of AI itself—from early perceptrons to advanced models like GPT. Ultimately, AGITB’s “all-tests-must-pass” philosophy ensures that AGI is evaluated according to the same fundamental principles that underlie biological intelligence.
While individual AGITB tests are trivial to solve in isolation, the true challenge lies in developing a universal AGI architecture capable of mastering all tasks within the unified framework. Since AGITB is a newly proposed benchmark, no classical imperative solution has yet emerged—nor can we rule out that one might. If a classical symbolic system were able to satisfy all AGITB conditions, it would serve as definitive proof that AGITB is inadequate as a test for genuine AGI.
History shows that hand-coded, task-specific systems struggle to scale toward general intelligence, highlighting the need for adaptive, learning-based architectures such as artificial neural networks. However, because state-of-the-art networks depend on pretraining with symbolic or structured data—which inevitably introduces bias and bypasses the grounding of meaning in raw signals—it remains an open and critical question how such models might begin learning directly from unstructured input, as required to pass the AGITB.